A ladder is resting against a wall. The top of the ladder touches the wall at a height of 12 ft. Find the length of the ladder if the length is 4 ft more than its distance from the wall. Answer should be in ft.
Answer by scott8148(6628)

by Pythagoras ___ 12^2 + (x - 4)^2 = x^2

144 + x^2 - 8x + 16 = x^2

160 = 8x